Country A has a Gini coefficient of .7 (point 7), and country B has a Gini coefficient of .4 (point 5). Which of the following can we conclude?

Country A has more income inequality than country B.
If you add the incomes of the populations of country A and country B, you will end up with a perfectly equal income distribution.
Both countries have equal income distributions, because their Gini coefficients are less than 1.
All of the above.

Which of the following represents an example of a regressive income tax?

Someone with an income of $20,000 pays $2,000 in taxes to the government, and a person with income of $30,000 pays $4,000 in taxes.
Someone with an income of $20,000 pays $2,000 in taxes to the government, and a person with income of $80,000 pays $7,000 in taxes.
Someone with an income of $20,000 pays $2,000 in taxes to the government, and a person with income of $40,000 pays $5,000 in taxes.
Someone with an income of $20,000 pays $2,000 in taxes to the government, and a person with income of $40,000 pays $4,000 in taxes.
Answer #1

1. Ans: Country A has more income inequality than country B.

Explanation:

Gini coefficient of zero means income is equally distributed whereas one means income is unequally distributed.

In the above statement , country A has a Gini coefficient of .7 which is closer to 1.

So country A has more income inequality than country B.

2. Ans: Someone with an income of $20,000 pays $2,000 in taxes to the government, and a person with income of $80,000 pays $7,000 in taxes.

Explanation:

Regressive income tax means decrease in tax rate when income increases.

Someone with an income of $20,000 pays $2,000 in taxes to the government. In this case,

tax rate = 2000 / 20000 = 10%

A person with income of $80,000 pays $7,000 in taxes.In this case ,

tax rate = 7000 / 80000 = 8.75 %